A 7.82 g sample of water is introduced into a 6.46 L flask containing some CO2 gas. The flask is heated to 307.11 °C at which temperature all of the water is converted to the gaseous phase, giving a total pressure in the flask of 12.305 atm. Calculate PH2O (in atm) in the flask at 307.11 °C. 


1
Expert's answer
2022-01-18T09:39:02-0500

P( h2o)= nRT/V

molar mass of water=18g/mol

total no. of moles of water=(7.82÷18)mol

=0.4344mol

but R=0.0821 L.atm/mol.k

T=(307.11+273) =580.11k

V= 6.46L

Therefore; pressure of water

=(0.4344×0.0821×580.11)÷6.46

= 3.203 atm